OpenStudy (anonymous):

Clarksburg, West Virginia, and Miami, Florida, lie approximately on the same meridian. Clarksburg has a latitude of 39.5 N and Miami, 25.5 N . Find the distance between these two cities. (The radius of the earth is 3960 mi.) Please round your answer to the nearest mile

OpenStudy (anonymous):

The distance between two points on a circle (or between two points on a sphere that have the same longitudinal coordinate) is given by the radius * the angle between them.

OpenStudy (anonymous):

Your angle needs to be in radians, however.

OpenStudy (anonymous):

Do you know how to convert degrees to radians?

OpenStudy (anonymous):

uhm x 180/pi

OpenStudy (anonymous):

Close. degrees *pi/180

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

802 miles

OpenStudy (anonymous):

I get something different. Let me check my numbers really fast.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Yeah, I've checked a few times. I think you have a minor mistake somewhere. 39.5 - 25.5 = 14. 14*3.14159265/180 = 0.244346095 0.244346095*3960 = 967.6

OpenStudy (anonymous):

oh i see my mistake haha, thanks for correcting me!:)